OpenStudy (anonymous):

Let u=<−3,3>, v=<1,−1>, and w=<−2,4>. Find the vector x that satisfies 4u−v+x=9x+w−. What is vector x

OpenStudy (april115):

vector?

OpenStudy (anonymous):

if u can't help its ok

OpenStudy (anonymous):

Check again if your equation is complete: \[ \Large 4u−v+x=9x+w−. \] What is the -. at the end?

OpenStudy (anonymous):

sorry that shouldn't be there

OpenStudy (anonymous):

So your equation is: \[\Large 4u−v+x=9x+w\]

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

Let u=<−3,3>, v=<1,−1>, and w=<−2,4> .

OpenStudy (anonymous):

So this equation is linear right? Solve for \(x\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

And that's your answer already.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[\Large x=\frac{1}{8}(4u-v-w) \]

OpenStudy (anonymous):

what does 4u-v-w=